# The Converse Madelung Question Schrödinger Equation from Minimal Axioms

### v1

@@ 4  Hamiltonian and Fisher Curvature
where $\alpha>0$ sets the regularisation scale. The first two terms reproduce classical mechanics; the last introduces Fisher curvature. Cross terms such as $\nabla\rho\!\cdot\!\nabla S$ are excluded by global phase symmetry (they reduce to a boundary divergence or select a preferred direction), and the prefactor is fixed by matching the continuity flux via the canonical bracket (Lemma 2.1). Any local scalar containing $\nabla S$, including $\nabla\rho\!\cdot\!\nabla S$ and $|\nabla S|\,|\nabla\rho|$, is either a total divergence or violates global $\mathrm{U}(1)$ or the $S$-shift generator role, and is excluded. A complete catalogue of first-derivative scalar candidates built from $\rho$ and $S$ is recorded in Appendix B, each tagged as divergence, $\mathrm{U}(1)$ violation, generator conflict, or admissible.
@@
Within the stated local first-order Hamiltonian class on $(\rho,S)$ and Axioms I-VI, the structural assumption is only that the regulariser be a local, positive, convex, first-derivative functional on $\rho$; no Fisher form is presupposed. We separate two routes: (i) an Euler-Lagrange uniqueness within the admissible class, and (ii) an operational projective-linearity stress-test after complexification. The present proposition establishes (i): analysis shows that $f(\rho)\propto 1/\rho$ is the sole admissible choice whose curvature term is time-reversal invariant and compatible with linear superposition once the unique local, derivative-free, gauge-covariant complexifier is imposed. Among all positive, rotationally invariant local quadratic functionals
@@
See Appendix D for full Euler-Lagrange and regularity conditions; complimentary superposition test supporting route (ii) is described in Appendix G.
@@
For Fisher-regularised dynamics one finds $\mathcal{R}_{\mathrm{proj}}(t)=0$ up to numerical tolerance, whereas any admissible departure from the Fisher form yields $\mathcal{R}_{\mathrm{proj}}(t)>0$ even under infinitesimal perturbations. Scripts are listed in the code archive (Appendix E); construction details are in Appendix G.
@@ 5  Hamilton-Jacobi System
These are the continuity and Fisher-regularised Hamilton-Jacobi equations. They form a closed, reversible system on $(\rho,S)$ with explicit time-symmetry invariance.
@@ 6  Emergence of the Schrödinger Equation
Write $R=\sqrt{\rho}$. Using (5.1)-(5.2) and
@@
For any admissible $f(\rho)\neq\kappa/\rho$ in the convex first-derivative regulariser, an $H^{-1}$-controlled state-dependent remainder persists in the Hamilton-Jacobi sector that cannot be cancelled by any local, gauge-covariant, derivative-free complexifier; consequently exact projective linearity fails. A concrete parameter identification is given in Appendix M; the functional-analytic no-go proof appears in Appendix D. For contrast, Doebner-Goldin diffusions lie outside our reversible Hamiltonian cone (breaking Axiom V), cf. [4, 5].
@@
Reversibility (Hamiltonian flow) and probabilistic composition select a projective complex representation of the dynamics (by Wigner’s theorem), so there must exist a complex structure in which time evolution is linear on rays. Throughout this section, linearisable means linearisable within Axioms I-VI, with a pointwise, derivative-free complexification on $(\rho,S)$. The local, gauge-covariant map $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ realises this structure in the hydrodynamic variables. Condition (6.2) is precisely what enforces exact projective linearity; it is therefore not an ansatz but the unique local complexifier within this class that renders the dynamics linear on $L^{2}$.
@@ 7.1 Gauge and Galilean invariance
$\mathbf{P}=\int\rho\,\nabla S\,dx\qquad(\text{so }\int\rho\,dx=1),$
@@
Indicating that the hydrodynamic representation already carries the projective representation of the Galilei group: mass enters as the central charge and need not be postulated independently.
In particular, $\{H,P\}=0$ for translation-invariant $V$, so $P$ is a conserved Noether charge. This realises precisely the Bargmann central extension of the Galilei algebra, with mass as the central charge ensuring the correct coadjoint-orbit structure [16], see Appendix I. The generator of boosts $\mathbf{K}=m\int\rho\,x\,dx-t\mathbf{P}$ satisfies
@@ 7.3 Dimensional analysis
Matching the free-particle dispersion fixes the scale: plane waves $\psi\sim e^{i(k\cdot x-\omega t)}$ in (6.3) obey $\omega=\hbar k^{2}/(2m)$, which combined with (6.2) gives $\alpha]=[\hbar^{2}/(2m)]$ and singles out the Fisher coefficient. Any other scaling fails to reproduce the quadratic dispersion mandated by Galilean kinematics.
@@ 8  Uniqueness of the Complexification
The local, pointwise, invertible, gauge covariant complexifier compatible with Axioms I-IV is $\psi=\sqrt{\rho}\,e^{iS/\kappa}$, which fixes $\alpha=\kappa^{2}/(2m)$. Any derivative-dependent or nonlocal map raises differential order and exits the class.
@@
Write $\psi=F(\rho)e^{iG}$ and compute the linear Schrödinger continuity law $\partial_{t}|\psi|^{2}+\nabla\!\cdot J=0$ with $J=\frac{\kappa}{m}\,\mathrm{Im}(\bar{\psi}\nabla\psi)=\frac{\kappa}{m}\,F(\rho)^{2}\,\nabla G.$ On the hydrodynamic side, (5.1) gives $\partial_{t}\rho+\nabla\!\cdot(\rho\,\nabla S/m)=0.$ Gauge covariance implies $G$ is affine in $S$ and independent of $\nabla S$, hence $G_{S}$ is a constant and $G_{\rho}$ is a scalar function. Matching the fluxes for arbitrary states forces
$\frac{\kappa}{m}\,F(\rho)^{2}\,G_{S}=\frac{\rho}{m}\quad\Rightarrow\quad G_{S}\equiv\frac{1}{\kappa},\ \ F(\rho)^{2}=\rho,$
@@
Equipping the $(\rho,S)$ phase space with the canonical symplectic form $\omega=\int d\rho\wedge dS$ and the Fisher metric in amplitude $g=\int 4\,|d\sqrt{\rho}|^{2}\,dx$ selects the integrable complex structure $J(d\sqrt{\rho})=\frac{1}{\kappa}\,dS$ compatible with $(\omega,g)$. Under the polar map $\psi=\sqrt{\rho}\,e^{iS/\kappa}$ the dynamics become linear on the projective Hilbert space (cf. [17, 18]).
@@ 9  Function-Space and Domain Considerations
We work on $\Omega=\mathbb{R}^{d}$ with decay $\rho,|\nabla S|\to 0$ as $|x|\to\infty$, or on smooth bounded $\Omega$ with boundary conditions that preserve integration by parts in the Hamiltonian. Consistent pairs are
$\text{Dirichlet: }R|_{\partial\Omega}=0,\ S|_{\partial\Omega}=\mathrm{const};\qquad\text{Neumann: }\nabla R\!\cdot n=0,\ \nabla S\!\cdot n=0.$
These preserve normalisation and energy conservation. On Lipschitz domains the standard trace theory justifies the integrations by parts used.
@@ 10  Topology and Vorticity
Recent developments have explored complementary routes linking Fisher information, hydrodynamics, and relativistic quantum theory. Fabbri [20] constructs a covariant “Madelung structure” for the Dirac equation, expressing spinor dynamics in polar variables as a coupled system of continuity, curl, and Hamilton-Jacobi-type equations built from first derivatives of the spinor fields. That approach is constructive: it reformulates an existing relativistic theory in hydrodynamic form. Our result is classificatory. Starting solely from locality, global phase and Euclidean invariance, reversibility, probability conservation, and convex regularity on $(\rho,S)$, we show that the canonical bracket and Fisher curvature together form the reversible information-hydrodynamic structure. The polar map $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ then emerges as the only admissible local lineariser, forcing the linear Schrödinger flow and fixing its scale. All such statements hold within the class defined by our axioms.
Information-theoretic works from a different direction reach compatible conclusions. Yang [21, 22] derives the Schrödinger and scalar field equations from an extended least-action principle that introduces vacuum fluctuations and information metrics, treating $\hbar$ as a minimal quantum of action and defining information curvature through relative entropy. These works show how information-based variational principles can reproduce and generalise the Fisher-regularised structure obtained here.
@@ 11  Many-Body and Spin Extensions
The following extension operates on configuration space $\mathbb{R}^{3N}$; $\rho(\mathbf{x}_{1},\!\dots,\!\mathbf{x}_{N})$ and $S(\mathbf{x}_{1},\!\dots,\!\mathbf{x}_{N})$ generate a formal hydrodynamics on this space. The term “hydrodynamic” here denotes the continuity and Hamilton-Jacobi structure rather than a literal fluid in physical three-space, consistent with the standard Madelung and Bohmian formulations. The scalar classification established earlier extends componentwise to multicomponent or spinorial fields. Extending to configuration space $\mathbb{R}^{3N}$, let $\rho(x_{1},\dots,x_{N},t)$ and $S(x_{1},\dots,x_{N},t)$ denote the single configuration-space density and phase, with $\nabla_{i}$ acting on $x_{i}$. Then
@@
so the probability currents are $j_{i}=\rho\,\nabla_{i}S/m_{i}$. With the polar map $\psi=\sqrt{\rho}\,e^{iS/\kappa}$ one has
$J_{i}=\frac{\kappa}{m_{i}}\,|\psi|^{2}\,\nabla_{i}(S/\kappa)=\rho\,\nabla_{i}S/m_{i}$
if and only if $\partial_{S}G\equiv 1/\kappa$ as in Proposition 8.1. The Madelung recombination on $\mathbb{R}^{3N}$ yields the linear $N$-body Schrödinger equation precisely when
@@
in precise analogy with the single-particle cancellation (cf. Eq. (6.2)), which we identify with a single Planck constant $\hbar=\kappa$ below.
@@
$\sum_{i}\alpha_{i}|\nabla_{i}\sqrt{\rho}|^{2}=\sum_{i}\frac{\hbar^{2}}{2m_{i}}\,|\nabla_{i}R|^{2}.$
@@
Sketch. A single local, gauge covariant complexifier on configuration space must be of the form $\psi=\sqrt{\rho}\,e^{iS/\kappa}$ with a constant $\kappa$ independent of $(x_{1},\dots,x_{N})$ by Proposition 8.1. Matching currents fixes $G_{S}\equiv 1/\kappa$, and recombination on each coordinate direction yields $\alpha_{i}=\kappa^{2}/(2m_{i})$. Hence a single $\kappa$ enforces a single Planck constant $\hbar=\kappa$ across all sectors.
@@
Let $\mathcal{H}_{N}$ be as above with masses $\{m_{i}\}$ and Fisher coefficients $\{\alpha_{i}\}$. If the reversible completion on configuration space admits a single local, gauge covariant complexifier $\psi=\sqrt{\rho}\,e^{iS/\kappa}$ that linearises the flow for all admissible data, then, for every $i$,
@@
Take factorised initial data $\rho=\prod_{i}\rho_{i}$, $S=\sum_{i}S_{i}(x_{i})$ with arbitrary one body pairs $(\rho_{i},S_{i})$. Linearity on $L^{2}$ under the same polar map requires the $i$th current to be $J_{i}=|\psi|^{2}\,\kappa^{-1}\nabla_{i}S$, which must match $j_{i}=\rho\,\nabla_{i}S/m_{i}$ for arbitrary $(\rho_{i},S_{i})$. Hence $\kappa$ is common and $G_{S}\equiv 1/\kappa$. Recombination along each coordinate gives the Laplacian quotient with coefficient $\alpha_{i}=\kappa^{2}/(2m_{i})$. Any attempt to use $\kappa_{i}$ depending on $i$ breaks exact projective linearity for superpositions mixing different masses. ∎
@@
The Hamiltonian acts on the full configuration-space wavefunction $\psi(\mathbf{x}_{1},\dots,\mathbf{x}_{N})$ without specifying symmetry. Bosonic and fermionic statistics enter as superselection conditions on the domain of $\psi$, not as modified dynamics: antisymmetry of $\psi$ automatically yields the effective Pauli pressure in marginal densities, while the underlying local Hamiltonian remains the same. Thus the same local Hamiltonian acts on bosonic or fermionic domains. Antisymmetry of $\psi$ yields the familiar effective Pauli pressure in reduced marginals without modifying the bracket or the Fisher curvature.

### v2

@@ 4  Hamiltonian and Fisher Curvature
where $\alpha>0$ sets the regularisation scale. The first two terms reproduce classical mechanics; the last introduces Fisher curvature. Cross terms such as $\nabla\rho\!\cdot\!\nabla S$ are excluded by reversibility and our first-order locality requirement: they are odd under $(t,S)\mapsto(-t,-S)$ and, when combined with the canonical bracket, generate higher-order contributions in the continuity equation that are not of the simple drift form $\nabla\!\cdot(\rho\nabla S/m)$.
Any local scalar containing $\nabla S$, including $\nabla\rho\!\cdot\!\nabla S$ and $|\nabla S|\,|\nabla\rho|$, is either a total divergence or conflicts with the $S$-shift generator role in Axiom II and the time-reversal symmetry of Axiom V. A complete catalogue of first-derivative scalar candidates built from $\rho$ and $S$ is recorded in Appendix B, each tagged as divergence, $\mathrm{U}(1)$ violation, generator conflict, or admissible.
@@
Within the stated local first-order Hamiltonian class on $(\rho,S)$ and Axioms I-VI, the structural assumption is only that the regulariser be a local, positive, convex, first-derivative functional on $\rho$; no Fisher form is presupposed.
We separate two routes: (i) an Euler-Lagrange uniqueness within the admissible class, and (ii) an operational projective-linearity stress-test after complexification.
The present proposition establishes (i): analysis shows that $f(\rho)\propto 1/\rho$ is the sole admissible choice whose curvature term is time-reversal invariant and compatible with linear superposition once the unique local, derivative-free, gauge-covariant complexifier is imposed. Among all positive, rotationally invariant local quadratic functionals
@@
See Appendix D for full Euler-Lagrange and regularity conditions; complementary superposition test supporting route (ii) is described in Appendix G.
@@
For Fisher-regularised dynamics one finds $\mathcal{R}_{\mathrm{proj}}(t)=0$ up to numerical tolerance, whereas any admissible departure from the Fisher form yields $\mathcal{R}_{\mathrm{proj}}(t)>0$ even under infinitesimal perturbations. Scripts are listed in the code archive Appendix E; construction details are in Appendix G.
@@ 5  Hamilton-Jacobi System
All subsequent uniqueness and linearisation statements are made for reversible completions of this system within the axiomatic class defined in Section 2. With the convention $Q_{\alpha}=-\alpha\,\Delta\sqrt{\rho}/\sqrt{\rho}$ fixed above, equation (5.2) can equivalently be written as $\partial_{t}S=-\frac{|\nabla S|^{2}}{2m}-V-Q_{\alpha}$, exposing the usual Bohm-Madelung structure.
@@ 6  Emergence of the Schrödinger Equation
Write $R=\sqrt{\rho}$. Using equation 5.2 and
@@
For any admissible $f(\rho)\neq\kappa/\rho$ in the convex first-derivative regulariser, an $H^{-1}$-controlled state-dependent remainder persists in the Hamilton-Jacobi sector that cannot be cancelled by any local, gauge-covariant, derivative-free complexifier; consequently exact projective linearity fails.
A concrete parameter identification is given in Appendix M; the functional-analytic no-go proof appears in Appendix D. For contrast, Doebner-Goldin diffusions lie outside our reversible Hamiltonian cone (breaking Axiom V) [4, 5].
@@
Assuming the reversible probabilistic theory admits a Hilbert-space realisation, reversibility (Hamiltonian flow) and probabilistic composition select a projective complex representation of the dynamics (by Wigner’s theorem), so there must exist a complex structure in which time evolution is linear on rays.
In the present work we use this only as a consistency criterion once a reversible probabilistic theory is given; it does not enter the hydrodynamic axioms themselves.
Throughout this section, linearisable means linearisable within Axioms I-VI, with a pointwise, derivative-free complexification on $(\rho,S)$. The local, gauge-covariant map $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ realises this structure in the hydrodynamic variables. Condition (6.2) is precisely what enforces exact projective linearity; it is therefore not an ansatz but the unique local complexifier within this class that renders the dynamics linear on $L^{2}$.
@@ 7.1 Gauge and Galilean invariance
$\mathbf{P}=\int\rho\,\nabla S\,dx\qquad(\text{so }\int\rho\,dx=1).$
@@
This indicates that the hydrodynamic representation already carries the projective representation of the Galilei group: mass enters as the central charge and need not be postulated independently.
In particular, $\{H,P\}=0$ for translation-invariant $V$, so $P$ is a conserved Noether charge. This realises precisely the Bargmann central extension of the Galilei algebra, with mass as the central charge ensuring the correct coadjoint-orbit structure [16], see Appendix I.
The generator of boosts $\mathbf{K}=m\int\rho\,x\,dx-t\mathbf{P}$ satisfies
@@ 7.3 Dimensional analysis
Matching the free-particle dispersion fixes the scale: plane waves $\psi\sim e^{i(k\cdot x-\omega t)}$ in (6.3) obey
$\omega=\frac{\hbar k^{2}}{2m},$
which combined with (6.2) gives
$\alpha]=\left[\frac{\hbar^{2}}{2m}\right]$
and singles out the Fisher coefficient. Any other scaling fails to reproduce the quadratic dispersion mandated by Galilean kinematics.
@@ 8  Uniqueness of the Complexification
The local, pointwise, invertible, gauge covariant complexifier compatible with our axioms is
$\psi=\sqrt{\rho}\,e^{iS/\kappa},$
which fixes
$\alpha=\frac{\kappa^{2}}{2m}.$
Any derivative-dependent or nonlocal map raises differential order and exits the class.
@@
Write $\psi=F(\rho)e^{iG}$ and compute the linear Schrödinger continuity law $\partial_{t}|\psi|^{2}+\nabla\!\cdot J=0$ with
$J=\frac{\kappa}{m}\,\mathrm{Im}(\bar{\psi}\nabla\psi)=\frac{\kappa}{m}\,F(\rho)^{2}\,\nabla G.$
On the hydrodynamic side, (5.1) gives $\partial_{t}\rho+\nabla\!\cdot(\rho\,\nabla S/m)=0.$ Gauge covariance implies $G$ is affine in $S$ and independent of $\nabla S$, hence $G_{S}$ is a constant and $G_{\rho}$ is a scalar function. Matching the fluxes for arbitrary states forces
$\frac{\kappa}{m}\,F(\rho)^{2}\,G_{S}=\frac{\rho}{m}\quad\Rightarrow\quad G_{S}\equiv\frac{1}{\kappa},\quad F(\rho)^{2}=\rho,$
@@
Equipping the $(\rho,S)$ phase space with the canonical symplectic form
$\omega=\int d\rho\wedge dS$
and the Fisher metric in amplitude
$g=\int 4\,|d\sqrt{\rho}|^{2}\,dx$
selects the integrable complex structure
$J(d\sqrt{\rho})=\frac{1}{\kappa}\,dS$
compatible with $(\omega,g)$. Under the polar map $\psi=\sqrt{\rho}\,e^{iS/\kappa}$ the dynamics become linear on the projective Hilbert space ([17, 18]).
@@ 10  Topology and Vorticity
Recent developments have explored complementary routes linking Fisher information, hydrodynamics, and relativistic quantum theory. Fabbri [20] constructs a covariant “Madelung structure” for the Dirac equation, expressing spinor dynamics in polar variables as a coupled system of continuity, curl, and Hamilton-Jacobi-type equations built from first derivatives of the spinor fields.
That approach is constructive: it reformulates an existing relativistic theory in hydrodynamic form. Our result is classificatory. Starting solely from locality, global phase and Euclidean invariance, reversibility, probability conservation, and convex regularity on $(\rho,S)$, we show that the canonical bracket and Fisher curvature together form the reversible information-hydrodynamic structure. The polar map $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ then emerges as the only admissible local lineariser, forcing the linear Schrödinger flow and fixing its scale. All such statements hold within the class defined by our axioms.
Information-theoretic works from a different direction reach compatible conclusions. Yang [21, 22] derives the Schrödinger and scalar field equations from an extended least-action principle that introduces vacuum fluctuations and information metrics, treating $\hbar$ as a minimal quantum of action and defining information curvature through relative entropy.
These works show how information-based variational principles can reproduce and generalise the Fisher-regularised structure obtained here. See also Yang’s fermionic field quantisation via an extended stationary action with a relative-entropy correction, which derives the Floreanini-Jackiw Schrödinger functional and verifies Poincaré algebra closure within the same information-geometric spirit.
@@ 11  Many-Body and Spin Extensions
The requirement that a single pointwise complex structure serves all subsystems forces a common information scale across different species, so separate particle types cannot carry independent adjustment knobs for the underlying curvature.
The following extension operates on configuration space $\mathbb{R}^{3N}$; $\rho(\mathbf{x}_{1},\!\dots,\!\mathbf{x}_{N})$ and $S(\mathbf{x}_{1},\!\dots,\!\mathbf{x}_{N})$ generate a formal hydrodynamics on this space. The term “hydrodynamic” here denotes the continuity and Hamilton-Jacobi structure rather than a literal fluid in physical three-space, consistent with the standard Madelung and Bohmian formulations.
The scalar classification established earlier extends componentwise to multicomponent or spinorial fields. Extending to configuration space $\mathbb{R}^{3N}$, let $\rho(x_{1},\dots,x_{N},t)$ and $S(x_{1},\dots,x_{N},t)$ denote the single configuration-space density and phase, with $\nabla_{i}$ acting on $x_{i}$. Then
$\mathcal{H}_{N}=\int\left[\sum_{i=1}^{N}\frac{\rho\,|\nabla_{i}S|^{2}}{2m_{i}}+V(\{x_{j}\})\,\rho+\sum_{i=1}^{N}\alpha_{i}\,|\nabla_{i}\sqrt{\rho}|^{2}\right]dx_{1}\cdots dx_{N}.$
@@
so the probability currents are $j_{i}=\rho\,\nabla_{i}S/m_{i}$.
With the polar map $\psi=\sqrt{\rho}\,e^{iS/\kappa}$ the configuration-space quantum current is
$J_{i}=\frac{\hbar}{m_{i}}\,\Im\bigl(\psi^{*}\nabla_{i}\psi\bigr),\qquad\hbar=\kappa,$
which coincides with the hydrodynamic current $j_{i}=\rho\,\nabla_{i}S/m_{i}$ for all admissible data if and only if $G_{S}\equiv\partial_{S}G\equiv 1/\kappa$, as in Proposition 8.1.
The Madelung recombination on $\mathbb{R}^{3N}$ yields the linear $N$-body Schrödinger equation precisely when
@@
in precise analogy with the single-particle cancellation (Eq. (6.2)), which we identify with a single Planck constant $\hbar=\kappa$ below.
@@
$\sum_{i}\alpha_{i}|\nabla_{i}\sqrt{\rho}|^{2}=\sum_{i}\alpha_{i}\,|\nabla_{i}R|^{2}=\sum_{i}\frac{\hbar^{2}}{2m_{i}}\,|\nabla_{i}R|^{2},$
where $\alpha_{i}=\hbar^{2}/(2m_{i})$ for each particle.
@@
Sketch. A single local, gauge covariant complexifier on configuration space must be of the form
$\psi=\sqrt{\rho}\,e^{iS/\kappa}$
with a constant $\kappa$ independent of $(x_{1},\dots,x_{N})$ by Proposition 8.1. Matching currents fixes $G_{S}\equiv 1/\kappa$, and recombination on each coordinate direction yields
$\alpha_{i}=\frac{\kappa^{2}}{2m_{i}}.$
Hence a single $\kappa$ enforces a single Planck constant $\hbar=\kappa$ across all sectors.
@@
Let $\mathcal{H}_{N}$ be as above with masses $\{m_{i}\}$ and Fisher coefficients $\{\alpha_{i}\}$. If the reversible completion on configuration space admits a single local, gauge covariant complexifier
$\psi=\sqrt{\rho}\,e^{iS/\kappa}$
that linearises the flow for all admissible data, then, for every $i$,
@@
Take factorised initial data $\rho=\prod_{i}\rho_{i}$, $S=\sum_{i}S_{i}(x_{i})$ with arbitrary one body pairs $(\rho_{i},S_{i})$. Under the Schrödinger evolution with parameter $\kappa$ the $i$th probability current is
$J_{i}=\frac{\kappa}{m_{i}}\,\Im(\psi^{*}\nabla_{i}\psi)=\frac{\rho}{m_{i}}\,\nabla_{i}S,$
so matching currents for arbitrary $(\rho_{i},S_{i})$ is automatic once $G_{S}\equiv 1/\kappa$ as in Proposition 8.1. Recombination along each coordinate gives the Laplacian quotient with coefficient $\alpha_{i}=\kappa^{2}/(2m_{i})$. Appendix L shows that any attempt to use $\kappa_{i}$ depending on $i$ breaks exact projective linearity for superpositions mixing different masses. ∎
@@
The Hamiltonian acts on the full configuration-space wavefunction $\psi(\mathbf{x}_{1},\dots,\mathbf{x}_{N})$ without specifying symmetry.
Bosonic and fermionic statistics enter as superselection conditions on the domain of $\psi$, not as modified dynamics: antisymmetry of $\psi$ automatically yields the familiar effective Pauli pressure in marginal densities, while the underlying local Hamiltonian, bracket, and Fisher curvature remain unchanged.